1. A parallel plate capacitor has an area of ​​1,000 cm ^ 2 and a distance between the plates of 1 cm. A dielectric material is inserted and its new capacitance is calculated to be 1.77 x 10 ^ -10 F. Determine the value of the dielectric constant.
a.

2

b.

2.25

c.

1.5

d.

0.75
Answer #1

2 Area of plates of capacitor = A=1000 cm 2 ( 1 cm = 10-2 m =) (1 cm)2 = (1.-2m) 2 1 cm ² 10-4 m²) 2 = 2 4 A = 1000 x 6 m A =